### v1

# BLAZE: Cross-Language and Cross-Project Bug Localization via Dynamic Chunking and Hard Example Learning

Abstract
Software bugs require developers to exert significant effort to identify and resolve them, often consuming about one-third of their time. Bug localization, the process of pinpointing the exact source code files that need modification, is crucial in reducing this effort. Existing bug localization tools, typically reliant on deep learning techniques, face limitations in cross-project applicability and effectiveness in multi-language environments.
Recent advancements with Large Language Models (LLMs) offer detailed representations for bug localization. However, they encounter challenges with limited context windows and mapping accuracy. To address these issues, we propose BLAZE, an approach that employs dynamic chunking and hard example learning. First, BLAZE dynamically segments source code to minimize continuity loss. Then, BLAZE fine-tunes a GPT-based model using challenging bug cases, in order to enhance cross-project and cross-language bug localization. To support the capability of BLAZE, we create the BeetleBox dataset, which comprises 26,321 bugs from 29 large and thriving open-source projects across five different programming languages (Java, C+++++ +, Python, Go, and JavaScript). Our evaluations of BLAZE on three benchmark datasets—BeetleBox, SWE-Bench, and Ye et al.—demonstrate substantial improvements compared to six state-of-the-art baselines. Specifically, BLAZE achieves up to an increase of 120% in Top 1 accuracy, 144% in Mean Average Precision (MAP), and 100% in Mean Reciprocal Rank (MRR). An extensive ablation study confirms the contributions of our pipeline components to the overall performance enhancement.
Index Terms:
Bug Localization, Embedding, Contrastive Learning
I Introduction
Software bugs are deviations from intended functionality in software applications, leading to undesired outcomes [1]. Resolving these bugs requires substantial developer effort, mainly to identify the exact source code file that needs correction. Prior research indicates that debugging consumes about one-third of developer work time [2, 3]. The challenge is exacerbated by the often overwhelming number of bug reports compared to the available developer resources, which can negatively affect customer satisfaction [4, 5].
Bug localization involves pinpointing the specific source code files that need modification based on the details provided in bug reports [6]. Effective bug localization tools minimize the time required to identify the root causes of defects by automating the tracing of errors to their sources [7].
Previous works (e.g.,  [8, 9, 10, 11, 12]) have introduced bug localization tools that measure the similarity between source code files and bug reports. These tools, which often employ deep learning techniques like Word2Vec or GloVe embeddings, are effective but heavily reliant on labeled data and generally require retraining for each new project, limiting their cross-project applicability [13, 14]. Although advancements have been made with tools that use transfer learning to function across different projects (e.g.,  [15, 16]), they still require labeled data from new projects for retraining. Moreover, these tools have predominantly been tested within single-language environments. This raises concerns about their effectiveness in cross-project bug localization involving multiple programming languages.
Recent studies have highlighted that Large Language Models (LLMs) provide detailed and context-rich representations for code retrieval tasks and bug diagnosis [17, 18]. This improved representation can enhance the ability of bug localization tools to accurately link bug reports with the corresponding source code files by integrating context-sensitive semantic understanding [19, 20]. However, such methods still encounter two major challenges. Firstly (Challenge 1: Limited Context Window), the typical context window of LLMs, which ranges from 512 to 2,048 tokens, is insufficient when compared to the average token length of 2,536 for a single file in common datasets [21].
Secondly (Challenge 2: Mapping Accuracy), such methods struggle to accurately map bug reports to the relevant buggy source code file, as illustrated in Figure 1(a), where the ideal non-overlapping distribution between buggy and non-buggy source code is not achieved. For this challenge, we highlight the need to prioritize the importance of source code files within projects. Consider the scenario depicted in Figure 2, where a font issue involves four files. The model erroneously associates the bug with ‘ThemeManager.java,’ an irrelevant file, and misses ‘FontController.java,’ which is crucial to the issue. Focused learning from such challenging examples can improve the model, aligning the representations of buggy source code files more closely with the bug report.
To overcome the above two challenges, we propose BLAZE, an approach that employs:
    Dynamic Chunking: Using dynamic programming, BLAZE segments source code at the boundaries of source code component declarations (e.g., class, interface, and method) to minimize continuity loss and adhere to context length constraints. We design these segments to be placed at component edges to prevent overlapping, resolving issues of continuity loss and excessive resource consumption caused by repetitive tokens.
    Hard Example Learning: BLAZE employs a GPT-based bug localization model tailored for identifying bugs across multiple languages and projects. We fine-tune it by prioritizing complex bugs (hard examples). BLAZE identifies such cases while tuning out less relevant outliers, which enhances the ability of BLAZE to generalize across different projects and languages and identifies critical misclassifications [22, 23]. As demonstrated in Figure 1(b), we find that the fine-tuning method could minimize the similarity overlap between buggy and non-buggy code files.
Refer to caption (a) Overlapped distribution before fine-tuning. Refer to caption (b) Minimally overlapped distribution after fine-tuning (BLAZE).
Figure 1: Kernel Density Estimate plot showing the distribution of similarities between buggy and non-buggy source code files and bug reports.
To evaluate the capability of BLAZE across languages, we compile BeetleBox, the most extensive cross-language and cross-project bug localization dataset to date, to the best of our knowledge. BeetleBox comprises 26,321 bugs that we mine from 29 real-world and thriving projects across five widely used programming languages: C++, Java, JavaScript, Python, and Go.
We evaluate the performance of BLAZE using three bug localization benchmarks, namely SWE-Bench [24], Ye et al. [21], and BeetleBox, which together encompass 31,745 bugs from 36 real-world software projects. BLAZE achieves up to 141% better Top 1 accuracy and 107.69% better MAP than three traditional cross-project bug localization tools [15, 5, 16]. Compared to modern language model-based tools [19, 20, 25], BLAZE achieves improvements of up to 63% in Top 1 and 69% in MAP. Moreover, we conduct an extensive ablation study, which confirms that our newly designed pipeline and fine-tuning techniques enhance performance. The results show that different components of our pipeline substantially contribute to improving bug localization performance to different degrees.
Contributions. In summary, this paper makes the following key contributions:
  • We introduce a novel bug localization technique (i.e., BLAZE). In this technique,
    • We propose a dynamic chunking technique to overcome the limitation of transformer-based approaches (limited context window).
    • We fine-tune BLAZE by learning from hard examples. BLAZE focuses on complex and frequently misclassified cases to enhance model performance in cross-language, cross-project settings.
  • We create BeetleBox, the largest to-date cross-language and cross-project bug localization dataset.
  • We conduct an ablation study to understand the importance of each component of the BLAZE pipeline.
Fostering open science. To foster future advances in bug localization, we release our dataset and we make the scripts used for our experiments publicly available [26].
Figure 2: Example of mining hard examples.
II Methodology
In this section, we describe our approach to designing BLAZE, which consists of three phases: Phase A introduces a fine-tuned language model to associate textual (bug report) and code elements (source code files). Phase B implements a dual-indexing system for source code files using both text and vector data from the fine-tuned model to facilitate a more effective retrieval of relevant files. Finally, in Phase C, we leverage the fine-tuned model from Phase A and the resulting indexed databases from Phase B together to facilitate the full process of matching incoming bug reports with the corresponding source code files. Below, we describe each phase in detail.
Phase A Fine-tuning of Language Model
Figure 3: Fine-tuning of language model (Phase A).
Figure 3 presents the steps we follow to fine-tune the model used in BLAZE. Each step of this process is detailed below.
Extraction of Source Code Elements. Our process begins with obtaining bug reports, the specific repository version where the bug exists, and all the source code files from that version provided in the targeted dataset (more details on how we collect and process the dataset in this study are described in Section III.B). In this initial step, we focus on extracting details such as the names and roles of components (e.g., methods, classes, and interfaces) from the collected source code files. Drawing on insights from prior research [27], which indicates that separate indexing of source code files enhances lexical retriever’s performance in code search tasks, we index source code files separately in the next phase of our pipeline (Phase B).
Moreover, we extract positional information, such as line numbers, which is essential for subsequent analysis stages. To identify all components within the source code, we use tree-sitter, [^1] https://github.com/tree-sitter/tree-sitter a tool that structures code into a graph format. From this tree-sitter graph, we extract all components along with their types, positions (line numbers), and names.
Input :
  cost_map: Map containing costs for each code component type.
  total_lines: Total number of lines.
  window_size: Maximum chunk size.
Output  :
  dp: Array containing minimum costs and corresponding breakpoints
1 Function dynamic_chunk(cost_map, total_lines, window_size)
2       line_map $\leftarrow$ get line number to component type map
3       split_cost_map $\leftarrow$ get line number to cost map
4       for $i\leftarrow$ 1 to total_lines do
5             minimum cost $\leftarrow$ $INFINITY$
6             breakpoint $\leftarrow$ $-1$
7             for $j\leftarrow$ MAX(i - window size, 0) to $i$ do
8                   cost = dp[$j$][0] + split_cost_map[$i$] or DEFAULT_COST
9                   if cost $<$ minimum_cost then
10                         minimum_cost = cost
11                         breakpoint = $j$
12
13
14            dp[$i$] = (minimum_cost, breakpoint)
15      return dp
16
Algorithm 1 Dynamic Chunking.
Dynamic chunking of source code file. In this step, we address the challenge of segmenting lengthy source code files into smaller segments that are suited for context-aware embedding models [28, 29, 30]. We employ a dynamic programming approach for chunking. This method segments files at natural boundaries (e.g., classes, interfaces, and methods), ensuring continuity and preventing overlaps—issues commonly associated with the traditional static chunking method and the sliding window technique.
Our algorithm, outlined in Algorithm 1, details the pseudo-code for our dynamic chunking approach. In the algorithm, we define ‘DEFAULT_COST’ as the cost of a random line. The approach begins by assigning each line of code a specific cost and categorizing it by component type. It then calculates the least expensive way to segment the code up to each line iteratively. It evaluates all potential breakpoints within the predefined window size (maximum context window supported by the intended model), updates to the minimum cost when a less expensive option is identified, and records the optimal breakpoint. The final output of the algorithm includes a list of the minimal costs associated with each breakpoint, alongside the breakpoints themselves. These are used to divide the source code files into smaller chunks.
Learning from hard examples. We fine-tune a Large Language Model (LLM), called CodeSage [31]. CodeSage is a GPT-based multi-modal embedding model. We adopt CodeSage to align bug reports and source code files within the same embedding space, drawing on its proven effectiveness in zero-shot code search tasks [31]. To optimize CodeSage for bug localization, we implement contrastive learning, which helps to distinguish between similar and dissimilar data pairs, i.e., this technique adjusts the model to minimize the distance between embeddings of bug reports and their corresponding buggy files (positive) while maximizing the separation from non-buggy files (negative). This distinction is guided by the Normalized Temperature-scaled Cross-Entropy (NT-Xent) loss, defined as:
$\mathcal{L}_{i,j}=-\log\frac{\exp^{sim(z_{i},z_{j})/\tau}}{\sum_{k=1}^{N}1_{[k\neq i]}\exp^{sim(z_{i},z_{k})/\tau}}$ (1)
where $z_{i}$ is the embedding of the bug report, $z_{j}$ is the embedding of the positive (buggy) source code file, and $\tau$ is the temperature parameter that scales the similarity. This loss function distinguishes between positive and negative examples by comparing similarity scores.
However, the standard NT-Xent loss may fall short in distinguishing between positive and negative source code files, potentially causing a bug report to align closer with a negative file. To mitigate this, we incorporate in-batch hard example mining into our training loop. This strategy prioritizes the most challenging cases: positive files with low similarity to the bug report and negative files with unexpectedly high similarity. We identify these hard examples by calculating the median similarity of all pairs in each batch. Negative source code files with similarity above this median and positive source code files below the median are considered hard examples. These cases are then given priority in the training process, with their impact on the loss function adjusted as follows:
$\begin{gathered}\mathcal{L}_{i,j}=-\log\frac{M_{i,j}*\exp^{sim(z_{i},z_{j})/\tau}}{\sum_{k=1}^{N}1_{[k\neq i]}\exp^{sim(z_{i},z_{k})/\tau}*M_{i,k}}\\ M_{m,n}=\begin{cases}\alpha,&\text{\parbox[t]{130.08731pt}{if $z_{n}$ negativeand $sim(z_{m},z_{n})$ $>$ $median(sim(z_{u},z_{v}))$}}\\ \beta,&\text{\parbox[t]{130.08731pt}{if $z_{n}$ positive and $sim(z_{m},z_{n})$ $<$ $median(sim(z_{u},z_{v}))$}}\\ 1,&\text{otherwise}\end{cases}\\ \alpha>1;\beta>1;1\leq u,v\leq N\end{gathered}$ (5)
In an average project, approximately 700 source code files exist, with only 1-2 directly related to a reported bug, as evidenced in Apache project dataset [21]. This rarity of relevant files presents a considerable challenge in bug localization. Our approach of hard example mining and scaling enhances the gradients from challenging cases and helps maintain a balanced learning environment, preventing the model from being overwhelmed by outliers. Moreover, the continuous nature of the scaling multiplier preserves the convergence properties of the NT-Xent loss. By employing median-based scaling, our method could improve the model learning from difficult examples, thereby enhancing its convergence. Henceforth, we refer to the fine-tuned language model as the BLAZE model.
Figure 4: Offline indexing phase of BLAZE (Phase B).
Phase B Offline Indexing
We introduce the offline indexing phase to enhance bug localization by proactively processing and indexing source code files. As mentioned earlier, repositories contain (on average) about 700 source code files, and without this phase, developers could face delays due to real-time data processing.
As depicted in Figure 4, upon receiving a bug report, BLAZE employs two distinct indexing strategies: the deep retriever stores data in the Milvus vector database [32], while the lexical retriever uses the no-SQL database ElasticSearch [33]. Having two types of retrievers can help to offset their weaknesses and take advantage of their strengths. Note that this phase builds upon the earlier phase of extracting names and positions of components (see Phase A). Below, we describe each retriever in detail.
Indexing for lexical retriever. Lexical retrievers are fast and use keyword matching to assess similarity. After extracting component names and positions as detailed in Phase A), each source code file is indexed in ElasticSearch. ElasticSearch leverages the Okapi BM25 algorithm [34] to compute similarity scores, a technique well-documented in previous studies for applications in code retrieval [27, 35]. Additionally, we index filenames and file paths to match logs and stack traces within bug reports.
Indexing for deep retriever. Deep retrievers integrate semantics to calculate the similarity between bug reports and source code files. Starting with the dynamically chunked source code files—as outlined in  Phase A—each chunk is transformed into an embedding using the BLAZE model. These embeddings are then indexed in Milvus, an open-source vector database specifically optimized for storing and retrieving embedding vectors in artificial intelligence applications.
Phase C Retrieval Pipeline
Figure 5: Retrieval phase of BLAZE (Phase C).
Figure 5 illustrates the retrieval phase of BLAZE. We leverage a combination of lexical and deep retrievers along with the fine-tuned BLAZE model. As shown in the figure, upon the arrival of a bug report, we query the Milvus and ElasticSearch indices to pinpoint the relevant source code files. We then integrate the rankings from both indices using a technique known as Reciprocal Rank Fusion (RRF). Next, we provide a detailed description of these steps.
Retrieval using deep retriever. Deep retrievers have demonstrated superior performance in ranking tasks compared to traditional vocabulary matching approaches [36]. To use a deep retriever, we first calculate the embedding of the bug report using our fine-tuned BLAZE model. Then, we retrieve the embedding of all the source code files from Milvus, a vector index we built during offline indexing. Milvus delivers the top K most similar chunks to the bug report embedding. We use cosine similarity to assess the similarity between embeddings. Finally, we sort the source code files based on decreasing similarity.
Retrieval using lexical retriever. In the lexical retrieval process, we query the ElasticSearch index with the bug report, which returns similarity scores for each source code file based on lexical matches. We use these scores to rank the source code files, leveraging ElasticSearch text searching capabilities to localize files potentially containing the bug. By directly using the bug report as a query, we ensure that the retrieval is tightly focused on the textual content relevant to the bug.
Reciprocal rank fusion. While deep retrievers demonstrate superior performance in capturing semantic similarity, they sometimes struggle with lengthy documents and may face challenges such as domain shifts when applied across different projects [37, 38]. These shifts can lead to outdated understandings of new domains. Conversely, lexical retrievers are immune to domain shifts but can encounter difficulties with vocabulary mismatches due to their lack of a semantic understanding. To address these limitations and harness the strengths of both retrieval types, we employ Reciprocal Rank Fusion (RRF) [39]. RRF integrates the rankings from both the lexical and deep retrievers by considering the position of each source code file in the rankings rather than merging their numerical scores. This method, supported by prior research [39, 40, 41], has proven to be a reliable and powerful technique for creating robust ensembles that enhance overall retrieval performance.
Given a bug report $r$, a source code file $s$, and a set of retrievers $M$, where $\pi^{m}(r,s)$ represents the ranking by retriever $m$ for source code file $s$ and bug report $r$, the RRF score is defined as follows:
$RRF(r,s,M)=\sum_{m\epsilon M}\frac{1}{k+\pi^{m}(r,s)}$ (6)
Here, $k$ is a hyperparameter. Based on prior studies [39, 41], $k=60$ has been found to provide the highest mean average precision (MAP). We also experimented with $k$ values ranging $\epsilon[50,70]$ in the BeetleBox dataset and confirmed that $k=60$ continues to yield the highest MAP.
III Experimental Design
In this section, we present the Research Questions (RQs) we tackle in our study, the process of our dataset collection, the subject baselines, and the performance metrics employed to evaluate the performance of BLAZE.
III.A Research Questions
Below, we present our RQs by explaining the motivation behind each one.
RQ1. How effective is BLAZE compared to state-of-the-art cross-project bug localization tools?
A plethora of bug localization tools have been proposed (e.g.,  [12, 42, 43, 16]), with cross-project tools (e.g., [15, 16]) being particularly effective. However, these tools still require project-specific additional training. Our proposed approach, BLAZE, is designed to function without this requirement. This RQ evaluates how well BLAZE performs compared to state-of-the-art cross-project bug localization tools.
RQ2. How effective is BLAZE compared to state-of-the-art embedding-based bug localization tools?
Embedding-based techniques, which assess document similarity, are extensively used for bug localization. These techniques differ from traditional bug localizers that provide binary predictions by producing embeddings and employing metrics like cosine distance to evaluate if a source code file contains bugs. Their ability to contextualize associations between bug reports and source code enhances their effectiveness. This RQ explores how well BLAZE performs compared to these embedding-based tools.
III.B Dataset Curation of BeetleBox
We have compiled an extensive multi-language, multi-project bug localization dataset. Previous research [21, 5, 44, 45] has proposed various datasets for bug localization, but these have numerous shortcomings, such as support for only one programming language and a limited range of projects. Additionally, some datasets, such as Bench4BL [46], contain inaccuracies in their ground truth data [47]. To address these challenges, we introduce BeetleBox. Below, we detail the process we follow for curating BeetleBox.
  • Selection of repositories. We use the GitHub API search feature to collect repository names. We search for repositories written in the top five programming languages, Java, Python, C+++++ +, JavaScript, and Go, as ranked in the StackOverflow Survey. [^2] https://survey.stackoverflow.co/2022/#technology-most-popular-technologies We select the top 10 repositories for each language, sorting them in descending order by their star ratings and most recent updates. This step guarantees that the repositories are actively contributed to and popular within the open-source community.
  • Linking pull-requests and issues. We gather ‘closed’ issues for each repository and their corresponding fixing Pull Request (PR). We identify these PRs by using predefined GitHub keywords. [^3] https://docs.github.com/en/issues/tracking-your-work-with-issues/linking-a-pull-request-to-an-issue
  • Filtration of issues. An issue may have several PRs, but typically only one is merged. To minimize errors in our ground truth, we filter out issues with the ‘duplicate’ tag and verify using GitHub API that each issue is linked to a single, merged PR. We initially started with 60 repositories. However, in certain projects, we found no issues meeting all the criteria. In the end, we are left with 29 repositories across five programming languages.
For each bug report in BeetleBox, the dataset, we include the bug status, repository name, repository URL, issue ID, a list of files modified during the fix, the title of the bug report, its body, URL link to the merged PR, URL link to the issue, SHA values of before and after the fix, the date and time the bug was reported, and the date and time of the fixing commit. Overall, BeetleBox comprises 26,321 bugs extracted from 29 projects. Detailed statistics about BeetleBox and a list of the projects can be found in the online Appendix [26].
Finally, to assess the accuracy of the ground truth in BeetleBox, we use the method outlined by Kim et al. [47]. Our manual analysis reveals that BeetleBox has an error rate of 0.06%, with a 95% confidence level.
III.C Evaluation Benchmarks and Baselines
Benchmarks. In addition to our BeetleBox dataset, we employ SWE-Bench [24], and Ye et al. [21]. Similar to BeetleBox, Ye et al. dataset is also recognized for its low error rate in its ground truth. While BeetleBox and Ye et al. are solely bug localization datasets, SWE-Bench is a recent dataset that is a benchmarking suite for bug localization as well as automatic patch generation.
TABLE I: Statistics on the used datasets, showing the distribution of bugs across training and testing, categorized by language for each dataset.
Dataset Purpose Language # Bugs
BeetleBox train c+++++ + 3,868
go 758
java 3,369
javascript 1,974
python 3,215
test c+++++ + 4,783
go 400
java 2,270
javascript 3,085
python 2,599
SWE-Bench [24] test python 2,294
Ye et al. [21] test java 16,314
In our methodology, we fine-tune the language model using the training split of BeetleBox. For evaluation, we use the test splits of BeetleBox, SWE-Bench, and the dataset from Ye et al. Table I provides brief statistics of these three datasets.
Studied baselines. We compare BLAZE with state-of-the-art cross-project bug localization tools, TRANP-CNN [15] and CooBa [16], which both use few-shot learning for cross-project adaptation. TRANP-CNN, a CNN-based deep learning tool, and CooBa, which uses dual encoders, train on a source project and are fine-tuned on a target project with limited samples. We also include BugLocator [5] as a baseline in our evaluation. Although not designed for cross-project use, BugLocator, like BLAZE, is an information retrieval-based tool that does not require additional training for deployment in new projects. The characteristics make it a relevant comparison in our evaluation.
In addition, we evaluate the performance of BLAZE compared to embedding-based bug localizers, namely SemanticCodeBERT [20] and FBL-BERT [19]. Both are BERT-based models, where SemanticCodeBERT incorporates semantic flow graph information into its training, and FBL-BERT incorporates changeset data. Recent advancements in generative AI have also introduced models like openai-embedding [25], which have demonstrated superior performance in tasks involving natural language and code. Therefore, OpenAI embedding is also included in our baseline for comparison.
III.D Evaluation Metrics
We adopt the following metrics that are widely adopted for evaluating bug localization tools [15, 43, 5]: Top N, Mean Reciprocal Rank (MRR), and Mean Average Precision (MAP).
  • Top N. Top N measures the overall ranking performance of the bug localization model. It indicates the percentage of bug reports for which at least one buggy source code file appears among the top N positions in the ranked list generated by the bug localization tool. Following previous studies [43, 15], we consider three values of N: 1, 5, and 10.
  • Mean Reciprocal Rank (MRR). MRR measures the average rank of the relevant file in the retrieved files set. It calculates the average reciprocal rank of the relevant source code files across all bug reports. The following equation calculates MRR, where $A$ represents the set of bug reports.
    $MRR=\frac{1}{|A|}\sum_{A}\frac{1}{Least\ rank\ of\ the\ relevant\ files}$ (7)
    For each bug report, our bug localization model ranks the source code files. Let us consider two bug reports, $report_{1}$ and $report_{2}$, each with six candidate source code files ranked by the model. For $report_{1}$, the ground truth of the retrieved files are $[0,0,1,0,1,0]$, and for $report_{2}$, the ground truth of the retrieved files are $[1,0,0,0,0,1]$. In this case, the least rank of relevant files is 3 and 1, respectively, for $report_{1}$ and $report_{2}$. Therefore, $MRR=\frac{1}{2}(\frac{1}{3}+\frac{1}{1})=0.67$.
  • Mean Average Precision (MAP). MAP measures the retrieval quality in scenarios where a bug is linked to multiple source code files. MAP considers the rank of all relevant files in the retrieved files list, making it a more comprehensive and unbiased metric than MRR. To compute MAP, we determine precision@1, precision@2, …, and precision@k; then, we calculate the average precision at different points. After calculating the average precision for each bug report, we calculate the mean of the average precision to compute MAP. The equation for calculating MAP is presented below, where $A$ represents the set of bug reports.
    $MAP=\frac{1}{|A|}\sum_{A}{AvgPrecision(Report_{i})}$ (8)
    To demonstrate the computation of MAP, consider the same example of two bug reports as previously discussed. The average precision calculated for $report_{1}$ is 0.37, while it is 0.67 for $report_{2}$. Thus, $MAP=\frac{1}{2}(0.36+0.67)=0.52$.
III.E Configuration Setup
We fine-tune the BLAZE model in a distributed setting with eight cores, 128GB of RAM, and fourteen Nvidia V100 32GB GPUs. We conduct inference on the same server using an Nvidia V100 32 GB GPU, which allows for fast and efficient predictions. We use PyTorch v.2.1.2, PyTorch-lightning v2.2.0, and the HuggingFace library v.4.36.2 to fine-tune our model. We also use ElasticSearch v8.7.1 and Milvus v2.3.4 for efficient similarity search. We fine-tune our language model on the BeetleBox dataset for 15 epochs, with batches of size 80 and a decaying learning rate starting at 2E-07, as recommended in prior work [28, 19]. More details about the hyperparameters are available in our online Appendix [26].
IV Results
In this section, we present our results per each RQ.
RQ1: Comparison to Cross-Project Tools.
TABLE II: Comparative performance of BLAZE and cross-project bug localization tools.
Dataset Model Name Top 1 Top 5 Top 10 MAP MRR
BeetleBox BLAZE 0.18 0.37 0.45 0.22 0.27
TRANP-CNN 0.08 0.13 0.19 0.09 0.13
CooBa 0.06 0.1 0.14 0.08 0.09
BugLocator 0.06 0.10 0.15 0.08 0.12
SWE-Bench BLAZE 0.29 0.68 0.79 0.43 0.45
TRANP-CNN 0.13 0.28 0.42 0.26 0.34
CooBa 0.12 0.25 0.38 0.21 0.3
BugLocator 0.1 0.14 0.21 0.17 0.21
Ye et al. BLAZE 0.16 0.35 0.44 0.21 0.26
TRANP-CNN 0.09 0.17 0.23 0.14 0.15
CooBa 0.08 0.13 0.19 0.09 0.13
BugLocator 0.09 0.11 0.16 0.07 0.12
Approach. We train our model using BeetleBox. We conduct the training on 13K bugs from five languages, covering 13 projects. Our evaluation dataset contains 31,745 bugs from five languages covering 29 real-world software projects. Following prior studies [24, 48], BeetleBox uses a completely different set of projects for training and testing splits. As previously presented in Table I, the SWE-Bench and Ye et al. datasets are used only for testing.
Result. Table II presents the performance of BLAZE compared to TRANP-CNN, CooBa, and BugLocator, across the studied datasets. Overall, BLAZE demonstrates its highest performance in the SWE-Bench dataset. Conversely, its performance is reduced in the Ye et al. dataset, as indicated by both Top 1 and MAP scores. Nonetheless, BLAZE consistently outperforms comparative baselines in this dataset and across all evaluated datasets.
In the BeetleBox dataset, BLAZE surpasses TRANP-CNN, with improvements ranging between 107% to 144.44%. Specifically, BLAZE achieves a 125% increase in Top 1 accuracy, a 144.44% increase in MAP, and a 107.69% in MRR. When compared to CooBa, BLAZE shows substantial gains: 200% in Top 1, 175% in MAP, and 200% in MRR. Against BugLocator, BLAZE shows improvements of 200% in Top 1, 175% in MAP, and 125% in MRR.
In the SWE-Bench dataset, BLAZE performs better than TRANP-CNN with increases between 32% and 123% across various metrics. The enhancements include a 123% improvement in Top 1, 65% in MAP, and 32% in MRR. Against CooBa, BLAZE shows a 141% increase in Top 1, 104% in MAP, and 50% in MRR. Compared to BugLocator, BLAZE shows improvements of 190% in Top 1, 52.9% in MAP, and 61.9% in MRR.
In the Ye et al. dataset, BLAZE shows an improvement over TRANP-CNN with gains of 77% in Top 1, 50% in MAP, and 73% in MRR. In comparison to CooBa, BLAZE shows increases of 100% in Top 1, 133% in MAP, and 100% in MRR. Against BugLocator, the model posts gains of 77.7% in Top 1, 200% in MAP, and 25% in MRR.
Additionally, across the studied programming languages, BLAZE achieves Top 1 scores of 0.18 in C+++++ +, 0.36 in Go, 0.26 in Java, 0.21 in JavaScript, and 0.3 in Python. These results show an improvement of 50.9% to 135.6% compared to TRANP-CNN, with the highest improvement observed in JavaScript and the lowest in Python. The trend is similar to that of CooBa, with improvements ranging from 82% to 162%, with the highest increase observed in C+++++ + and the lowest in Python. Against BugLocator, the improvement ranges from 87% to 160%, with the highest improvement in JavaScript and the lowest in Python.
RQ2: Comparison to Embedding-Based Tools.
TABLE III: Comparative performance of BLAZE and embedding-based bug localization tools.
Dataset Model Name Top 1 Top 5 Top 10 MAP MRR
BeetleBox BLAZE 0.18 0.37 0.45 0.22 0.27
OpenAI 0.13 0.26 0.32 0.1 0.19
SemanticCodeBERT 0.11 0.20 0.31 0.13 0.21
FBL-BERT 0.12 0.23 0.30 0.16 0.19
SWE-Bench BLAZE 0.29 0.68 0.79 0.43 0.45
OpenAI 0.27 0.63 0.72 0.36 0.39
SemanticCodeBERT 0.23 0.5 0.64 0.31 0.33
FBL-BERT 0.20 0.57 0.51 0.31 0.30
Ye et al. BLAZE 0.16 0.35 0.44 0.21 0.26
OpenAI 0.17 0.36 0.43 0.19 0.25
SemanticCodeBERT 0.13 0.24 0.30 0.17 0.22
FBL-BERT 0.10 0.24 0.28 0.14 0.18
Approach. We examine three embedding-based bug localization tools: SemanticCodeBERT [20], FBL-BERT [19], and OpenAI [25]. For SemanticCodeBERT [^4] https://github.com/IBM/semanticflowgraph and FBL-BERT [^5] https://anonymous.4open.science/r/fbl-bert-700C/README.md, we use their publicly available replication package to train the models and produce embeddings. For OpenAI, we obtain embeddings for both bug reports and source code files using the OpenAI API. [^6] https://openai.com/index/openai-api/ We employ cosine similarity to rank the source code files based on their similarity scores, facilitating the identification of the buggy files.
Result. Table III illustrates the performance of BLAZE relative to the studied embedding-based tools. From the table, we observe trends similar to those observed in Table II. We also observe that the performance across all models is strongest in the SWE-Bench dataset, which comprises Python projects, and weakest in the Ye et al. dataset, featuring Java projects.
In the BeetleBox dataset., BLAZE demonstrates superior performance to OpenAI, with improvements ranging from 38% in Top 1 accuracy to 120% in MAP and 42.11% in MRR. Similarly, when compared to SemanticCodeBERT, BLAZE achieves substantial gains, with increases of 63% in Top 1, 69% in MAP, and 28% in MRR. It also shows notable enhancements against FBL-BERT, including a 50% rise in Top 1 and improvements in MAP and MRR by 37% and 42%, respectively.
The performance trend continues in the SWE-Bench dataset, where BLAZE outperfomrs OpenAI with gains from 7% in Top 1 to 19% in MAP and 15% in MRR. It also outperforms SemanticCodeBERT with a 26% increase in Top 1 and even greater enhancements in MAP and MRR. Against FBL-BERT, BLAZE shows substantial improvements of 45% in Top 1, 38% in MAP, and 50% in MRR.
In the Ye et al. dataset, BLAZE slightly lags behind OpenAI in Top 1 by a margin of 5%, but it still achieves better performance in MAP and MRR by 10% and 4%, respectively. It demonstrates enhancements over SemanticCodeBERT, improving by 23% in both Top 1 and MAP, and 18% in MRR. Similarly, against FBL-BERT, BLAZE achieves substantial performance increases of 60% in Top 1, 50% in MAP, and 44% in MRR.
When analyzing the performance of BLAZE across the studied programming languages, we find that BLAZE consistently outperforms the studied models in Top-1 scores across most programming languages. The exception is Go, where OpenAI slightly outperforms BLAZE in Go projects. We find that Go projects contain only 758 bugs, the fewest in our dataset. Additionally, the unique concurrency model and error-handling patterns of Go may pose challenges for BLAZE and contribute to lower performance in this language. Moreover, BLAZE surpasses OpenAI with improvements ranging from 14.24% to 16.86% in all languages except Go, where it lags by 0.66%. Compared to OpenAI, BLAZE achieves the highest improvement in Python and the lowest improvement in C+++++ +. Compared to SemanticCodeBERT, BLAZE shows consistent improvement between 15% and 41%, with the highest being observed in JavaScript and the lowest in Go. The trend continues when compared with FBL-BERT, where improvements range from 34.1% to 67.5%, peaking in Java and being lowest in JavaScript.
V Ablation Study
This section presents an ablation study to assess the influence of individual components in the BLAZE pipeline on its overall performance. Key strategies contributing to the effectiveness of BLAZE include (1) a combined pipeline of lexical and deep retrievers, (2) learning from hard examples, and (3) dynamic chunking. Next, we detail the results of the ablation study to evaluate the impact of each component.
V.A Impact of hybrid retrieval technique
Figure 6: Precision recall curves of the retrievers.
TABLE IV: Performance comparison between Lexical, Deep, and Hybrid Retriever.
Retriver Type Top 1 Top 5 Top 10 MAP MRR
Lexical 0.07 0.12 0.19 0.11 0.13
Deep 0.12 0.25 0.31 0.14 0.18
Hybrid (BLAZE) 0.18 0.37 0.45 0.22 0.27
In BLAZE, we use a hybrid technique by combining outputs from two types of retrievers: lexical and deep. The lexical retriever is typically useful at finding exact matches and managing simple searches, while the deep retriever is better suited for understanding complex queries through their semantics. We assess their combined performance with a precision-recall graph, as shown in Figure 6. The figure demonstrates how accurately BLAZE identifies buggy source code files at various decision thresholds, with higher precision and recall indicating better performance.
Figure 6 shows that the lexical retriever initially has high precision but low recall, indicating it identifies relevant files well but may miss others. As it retrieves more files, precision drops. Conversely, the deep retriever’s precision increases as it processes more files, particularly at higher recall levels. By combining both retrievers in the BLAZE pipeline (as denoted ‘Hybrid’ in Figure 6), we achieve a balance that enhances overall results.
Furthermore, we analyze the performance of each retriever and their combined effect in the BLAZE pipeline on the BeetleBox dataset. From Table IV, we observe that the deep retriever outperforms the lexical retriever, showing improvements of 71% in Top 1, 27% in MAP, and 38% in MRR. Combining the retrievers further enhances performance, showing improvements of 45% to 57% over the deep retriever alone. This result demonstrates that combining both retriever types mitigates their individual weaknesses and enhances the overall performance of BLAZE.
V.B Impact of fine-tuning
(a) Epoch 0 (CodeSage)
(b) Epoch 5
(c) Epoch 15 (BLAZE)
Figure 7: Scatter plots showing the chunks of the buggy and non-buggy files associated with the Apache Dubbo-1216 bug report. denotes bug report, denotes chunks from buggy files, and denotes chunks from non-buggy files.
In BLAZE, we fine-tune a GPT-based model [31] to better target hard-to-localize bugs by adjusting how closely buggy (positive files) and non-buggy (negative files) source code relate to bug reports (see Phase A). We assess the impact of this fine-tuning by comparing the performance of the model at different stages: epoch 0 (CodeSage), epoch 5, and epoch 15 (BLAZE). We use t-SNE [49], a method that simplifies complex data into two or three dimensions for easier visualization, to generate and display embeddings for a specific bug report (i.e., Apache Dubbo-1216) and it is corresponding source code.
Figure 7 shows that the ability of the CodeSage model to distinguish between buggy and non-buggy files in the analyzed bug report (i.e., Apache Dubbo-1216) evolves over time. At the start (epoch 0), the model struggles to differentiate, with overlapping positive and negative chunks. By epoch 5, there is a clear improvement: positive chunks start clustering more closely, and negative chunks spread out, although some overlap remains. By epoch 15, the model has made substantial progress; positive chunks are well-grouped, and no negative chunks are closer than the positives, indicating successful fine-tuning.
TABLE V: Performance of CodeSage vs. BLAZE.
Retriver Type Top 1 Top 5 Top 10 MAP MRR
CodeSageDeep 0.11 0.21 0.28 0.12 0.16
CodeSageHybrid 0.17 0.32 0.42 0.19 0.24
BLAZE Deep 0.12 0.25 0.31 0.14 0.18
BLAZE Hybrid 0.18 0.37 0.45 0.22 0.27
Also, we evaluate the CodeSage model on the BeetleBox dataset. Table V shows the results of the CodeSage and BLAZE model with different types of retrievers. For instance, CodeSageDeep reflects the performance with a deep retriever, while CodeSageHybrid indicates the results using a hybrid retriever combining deep and lexical retriever. Using only the deep retriever, BLAZE achieves a performance improvement of 6.9%–12.5% over CodeSage, including 9.1% for Top1, 16.7% for MAP, and 12.5% for MRR. The hybrid technique used by the BLAZE pipeline leads to even more enhancements, ranging from 5.8% to 15.7%, with gains of 5.8% in Top1, 15.8% in MAP, and 12.5% in MRR. These results support that fine-tuning the model on challenging (hard) examples improves its effectiveness.
V.C Impact of dynamic chunking
In BLAZE, we use dynamic chunking to manage the issue of limited context size while reducing continuity loss. Previous research [50, 51] has used methods like sliding windows and static chunking. The sliding window method overlaps chunks to maintain continuity but complicates similarity calculations due to these overlaps. Static chunking, on the other hand, avoids overlaps but can lead to a loss of continuity. Our approach uses dynamic programming to find breakpoints that reduce continuity loss without shrinking the window size. We trained and evaluated BLAZE using all three chunking methods to see the effects of each.
TABLE VI: Static vs. Sliding window vs. Dynamic chunking.
Model Top 1 Top 5 Top 10 MAP MRR
BLAZE Static + Deep 0.11 0.22 0.26 0.13 0.17
BLAZE Sliding + Deep 0.11 0.24 0.3 0.13 0.17
BLAZE Dynamic + Deep 0.12 0.25 0.31 0.14 0.18
BLAZE Static + Hybrid 0.15 0.33 0.37 0.18 0.25
BLAZE Sliding + Hybrid 0.17 0.36 0.42 0.21 0.25
BLAZE Dynamic + Hybrid 0.18 0.37 0.45 0.22 0.27
Table VI presents a performance comparison of various chunking techniques. For instance, BLAZEStatic + Deep represents the performance of static chunking paired with the deep retriever. The table shows that BLAZEDynamic + Deep improves performance by 5.5–16% over static chunking and by 3.23–8.33% over the sliding window method. Within the hybrid technique of BLAZE pipeline, dynamic chunking shows a performance increase of 7–18% compared to static chunking and 2–7.41% compared to sliding window.
VI Threats to Validity
Below, we discuss the threats to the validity of our study.
Internal validity. One threat to internal validity is the hyperparameters we used to train the models. To mitigate this, we use the same parameters used by several prior studies that use large language models [28, 19, 52, 13].
External validity. A threat to the external validity of our study lies in the generalizability of the results. We assess cross-project capabilities using separate projects for training and testing. Although the choice of test projects may affect performance, this approach aligns with previous studies that evaluate zero-shot and cross-project capabilities [48, 24]. Additionally, our evaluation of 31,745 real-world bugs ensures that our findings hold, even with different project selections, due to the large scale of our dataset. To address potential selection bias in the BeetleBox dataset, we collected project repositories based on star ratings and recent updates, ensuring that the repositories are actively contributed to and popular within the open-source community. This approach aims to include high-quality, widely used projects across the selected top five programming languages.
Construct validity. Our selected evaluation metrics can threaten the construct validity of the study. We used Top N, MAP, and MRR as evaluation metrics. These are well-known evaluation measures adopted in prior studies which are widely used when evaluating the performance of bug localization tools that rely on information retrieval techniques [19, 53, 21, 5].
VII Related Works
In this section, we present prior work related to bug localization approaches and discuss how our approach differs from them.
A plethora of work proposed lexical retriever-based approaches to localizing bugs. Such approaches use BM25 and TF-IDF techniques to measure the similarity between a bug report and source code files [5, 6, 54, 55, 56, 57]. For example, BugLocator [5] used TF-IDF to measure the similarity between the bug report and source code files using a revised vector space model. Sisman et al. [58] and Amalgam [6] used version history along with the VSM model to localize bugs. Shi et al. [59] found that BM25 and its variants are better than TF-IDF, VSM, and rVSM in the bug localization task. BoostNSift [55] used BM25 for method-level bug localization. Our approach, BLAZE, differs from prior work since it leverages the advantages of both the BM25-based retrieval method and deep learning-based technique to perform cross-project bug localization.
Other work proposed deep learning-based approaches to localizing bugs [53, 43, 5, 60, 11]. For example, FLIM [53] used CodeBERT embedding for calculating the similarity between bug reports and source code files at the function level. DNNLoc [61] combined a deep neural network with the VSM model to increase effectiveness across different types of metrics such as textual similarity, class name similarity, bug-fixing recency, and frequency. NP-CNN [43] employed a Convolutional Neural Network (CNN) to leverage both lexical and semantical properties. BL-GAN [12] proposed a generative adversarial network-based solution for bug localization. Several studies [42, 9, 8, 10] incorporated program structure information in the model to generate more useful program representation. For example, Liang et al. [10] and Li et al. [8] used abstract syntax trees, data flow graphs, and program flow graphs to generate a better representation of source code. While BLAZE employs an embedding model similar to the techniques discussed previously, it differs by its ability to be applied to new, unseen projects without requiring additional training.
Several studies offered test case coverage or spectrum-based solutions for bug localization [62, 63, 64, 65, 66, 67, 68, 69]. For example, FusionLocalizer [69] combines multiple test case-based fault localizers. However, their approach depends on tests, whereas BLAZE uses an information retrieval approach solely based on bug reports. Moreover, the selection step in the FusionLocalizer technique may introduce additional bias, while BLAZE employs Reciprocal Rank Fusion (RRF) to consider the ranks of all retrievers, effectively reducing bias. Similarly, UniLoc [70] uses information retrieval for localizing bugs in continuous integration tools, while BLAZE focuses on bug reports. Additionally, CrossLocator [71] introduces a translation-based method for projects that uses multiple languages for bug reports (e.g., English, Chinese), whereas BLAZE identifies bugs in projects involving different programming languages.
While numerous studies have been conducted on bug localization tools, most of them are project-context approaches, i.e., the research on cross-project bug localization is limited in comparison. For example, TRANP-CNN [15] used a combination of transfer learning and CNN models to attain competitive performance in cross-project bug localization at the file level. CooBa [16] follows a similar approach but employs a shared encoder to capture cross-project features. Such approaches still require further training before deploying and making them work for new projects. In contrast, BLAZE is designed in a way that it can work on a new project without further training. DSSDPP [72] investigated a sampling approach to address disparities between source and target projects, whereas xLoc [73] used transformers to identify bugs in cross-language API calls. Nevertheless, both tools concentrate on method-level defect prediction, in contrast to BLAZE, which localizes bugs at the file level.
Previous research has also explored embedding-based methods for bug localization. SemanticCodeBERT [20] integrates semantic flow graph data, while FBL-BERT [19] uses changeset data. However, neither focuses on hard examples to improve feature separation. In contrast, BLAZE learns from hard examples to significantly enhance model performance.
VIII Conclusion
In this paper, we introduce BLAZE, a bug localization tool that operates without project-specific training, facilitating cross-project and cross-language bug localization. BLAZE employs a dynamic chunking and hard example learning methodology to model lengthy source code files. To support cross-project and cross-language bug localizers, we curate BeetleBox—one of the largest (if not the largest) cross-project and cross-language bug localization datasets. We evaluate BLAZE with 31,745 bug reports from 34 real-world projects, showing up to a 100% improvement in the Top 1 metric against cross-project localizers and up to a 60% improvement over modern large language model approaches. Our results suggest that the use of dynamic chunking and hard example learning is key in enhancing the performance of bug localization tools.